Sol Summary – December 5th

Crew 184 Sol Summary Report  05 Dec 2017

MDRS Sol Summary Report for Sol 4

Summary Title:  Our first failure and major replan

Mission Status:  Today was a hectic day with three major failures our
team dealt with.
1.    Internet – Between 8:30am and 9:30am 800MB of bandwidth was
utilized. It is unknown what caused this but likely culprit was a
device being left on and doing a software update in the background.
Internet communication was lost at this time.  These reports are being
delivered by Shannon today.
2.    Water – During our morning fill of the loft tank the pump
was non-functional.  Possible causes were quickly narrowed down to
either a frozen pipe due to the below freezing temperatures or a
clogged filter. Our team immediately switched to using bottled water
which we keep as an emergency reserve while we troubleshot the issue.
At 1pm we called Mission Control for advice who recommended we connect
the pipe heating system and wait till midday, this resolved the issue
around 3pm and water was restored to the HAB
3.    Power – During our morning check of SOC it was noted that
the generator was not running, having failed sometime during the
night, and that SOC was at 3%.  We immediately instituted a shutdown
of MDRS powered equipment to preserve what little power we had and
investigated the generator.  It was noted that a significant oil leak
had developed from the oil cap, upon tightening the cap the leak
stopped and the generator was restarted.  When returning to the
science dome it was observed that SOC was hovering around 10% and did
not appear to be charging.  After consultation with Mission Control it
was discovered that the generator was in idle due to the power system
not requesting it to go to full delivered power given the fact that
the solar array system was charging the battiers (however slowly) .
In addition the SOC was likely incorrect due to it being a calculated
value and us confusing it due to going off scale.  We will assess SOC
tonight after a full day in the sun and decide forward plan.

Our morning was mostly consumed with maintenance and troubleshooting,
we quickly discovered that when receiving failures our level of
knowledge of the systems was exhausted in how to troubleshoot,
recommend future crew get access to and study manuals for MDRS
equipment prior to stay.

In the afternoon when we got our facilities in a stable configuratrion
four members of our team went on our first Matryoshka EVA. We were
able to go to the Kissing Camel ridge and collect two samples, however
we were unable to go to the other site in URC South due to delays in
getting out since we had to fix our non-nominal systems.

The two sample we collected were to highlight interesting
geomorphological features to be used in outreach activities back in
Oxford. The strike and dip of the ridge component the first sample was
extracted from was also taken. The intention of sample collection and
strike and dip measurement is to determine whether satellite and
outcrop scale imagery data is reliable at predicting these.

Sol Activity Summary:
1.    Troubleshooting and EVA as discussed above
2.    Exercise
3.    Assembling of exercise equipment using martian dirt
(medicine ball complete!)
4.    HAB Cleaning
5.    Solar observation

Look Ahead Plan:
With internet down we are unable to contact our scheduling team.  We
will receive the schedule in the morning when internet bandwidth
restarts and go over it as a team to plan our day.

Anomalies in work:
1.    Yamaha 250 ATV #3 wheel deemed non-recoverable.  We have
parked it off to the side for mission support.  Photos of damage
attached.

Weather: Very cold!  It was below freezing for most of the day.

Crew Physical Status: All in good health.

EVA: Delayed due to non-nominal systems, so we are requesting the same
EVA from Sol 3 but just for site A (mound on URC South) to finish off
the sample collection. The plan and crew utilised will be the exact
same as today.

Reports to be file:
Greenhab
Operations
Astronomy
EVA Request
Journalist
Sol Summary

Support Requested:
Trash placed in the Rear Airlock for pickup.  Burnable trash on the
side with the box.  Non-burnable trash on the other side.

Operations Report – December 5th

Crew 184 Operations Report 12/05/2017

SOL: SOL 4

Name of person filing report: Akash Trivedi

Non-nominal systems: Water (pump/lines – static to loft), generator
failure, sudden loss of data

Notes on non-nominal systems:  Water: tanks from static to loft not
transferring due to speculated frozen pipe or blockage. Waited until
10am as per protocol to see if frozen pipes thawed. At 10am, still
frozen. Waited until 12pm for sunlight to be in contact with pipes to
see if thaws. After the meeting at lunch, we learned that a plug by
the static tank needed to be seated. This allows heating along the
pipes. Once plugged we waited some time, and pump was operational
again at around 3pm. Anomaly resolved.

Generator: Noticed generator off at 0900 having failed during the
night. Turned on and seemed to be operating nominally. Returned to
check up on it at 1000, and noticed a coolant leak from the radiator.
Generator turned off and leak corrected by tightening the antifreeze
cap. Generator turned back on and checked status on power console –
seems to not be doing the charging as it’s on “absorb” and SOC not
increasing despite minimal hab power usage. After Lunch meeting, we
monitored thegenerator closely and at sunset the SOC returned to 100%.
Will continue to monitor closely, but power is good at this time.

Loss of internet data: Data check was conducted at 0845 when it was
approximately 800MB, and an hour later, this dropped to zero. Reason
unknown.

Generator (hours run):  Unknown due to night time failure

Solar— SOC % (Before generator is run at night) 4% at 0900, 14% at
1000 (when coolant leak discovered), static 9% on “absorb charging”
with minimal hab usage and generator powered on

Diesel – 48%

Propane –  79 percent volume

Ethanol Free Gasoline (5 Gallon containers for ATV) – 1.25 gallons

Water (trailer) – 600 gallons

Water (static) – 450 gallons

Trailer to Static Pump used – No

Water (loft) – Static to Loft Pump used – Yes

Water Meter: 46 gallons

Toilet tank emptied: Yes

ATV’s Used: Honda ATV, Yamaha 1

Oil Added? No

ATV Fuel Used: 1.25 gallons

# Hours the ATVs were Used today: 00:45

Notes on ATVs: Yamaha 3 Rear right tire is still wobbly, EVA team
assessed and determined it will require. Photos will be provided in
overall Sol 4 package for awareness and review.

Deimos rover used: Used

Hours: 00:45

Beginning charge: 100%

Ending charge: 100%

Currently charging:Yes

Sojourner rover used:  ASSIGNED TO DIRECTOR

Hours:

Beginning charge:

Ending charge:

Currently charging:

Spirit rover used:  Not used

Hours:

Beginning charge:

Ending charge:

Currently charging:

Opportunity rover used:  Not Used

Hours:

Beginning charge:

Ending charge:

Currently charging:

Curiosity rover used:  Not Used

Hours:

Beginning charge:

Ending charge:

Currently charging:

HabCar used and why, where?  No

General notes and comments: Multiple non-nominal systems!!!

Summary of internet: Sudden anomalous data loss – see above

Summary of suits and radios: All nominal based on last usage

Summary of Hab operations: Loft water tank almost empty, multiple
non-nominal systems

Summary of GreenHab operations:  All nominal

Summary of ScienceDome operations: All nominal

Summary of RAMM operations: Not Operational

Summary of health and safety issues: No problems now that water, heat,
and power are back.

Questions, concerns and requests to Mission Support: See above on
non-nominal systems. Lack of training and relevant manuals for
troubleshooting significant issues in mission critical systems. Poor
insulation of static tank to loft tank pipes – could lead to pipe
cracking

GreenHab Report – December 5th

GreenHab Report

Crewmember’s name: Akash Trivedi

Date: 5/12/17

Environmental control:

Ambient (no heating or/cooling): No wind, door open at 9am. Since
generator failed and heater would not work, we closed door to
stabalise temperature.

Heating: Nominal, once generator was back on

Cooling: Not operational

Both heating and cooling

Shade cloth on

Average temperature: 21 oC in morning before power shortage, 16 oC at 1700

Changes to crops:  Same tomato plant droopy, also noticed a droopy
radish plant and moved it to a brighter location

Daily water usage for crops: Nominal

Time(s) of watering for crops: 9am

Research observations: (Currently not operational)

Changes to research plants: N/A

Daily watering and amount of water used: 2 gallons

Aquaponics: (Currently not operational)

Narrative: Plants are doing good on the whole!

Support/supplies needed: Fire extinguisher required, extra mineral
salts for supplements in case of deficiencies

Journalist Report – December 5th

Crew 184 Journalist Report

Willie Schumann

05 December 2017

Title                            Crisis Management

Narrative                  Some people say you only really arrived to a new home, when you managed your first crisis and you didn’t run away. Well, that means we have finally arrived on Mars. We woke up to a bunch of problems, which could have had a dramatic effect on the livelihood of crew number 184.

Crew Engineer Hunt woke up early today and was the first to realize, that our internal water tank in our top floor had reached a critical low level. This tank fuels our kitchen, the bath and the toilet – it is the heart of our Martian shelter. It is just below the roof, because it uses gravity to deliver water to the various outlets below. At first we thought, our pump was defective, that enables the whole operation.

That conclusion was plausible, because shortly after crisis number one we detected crisis number two. Our battery, which is fueled by a generator and the solar panels at their respective working hours, was down to five percent. Commander Horn and Crew Engineer Hunt, still in his pajamas, observed the devices and detected a leaking oil tank.

Was our failing energy source the reason for the block in the water delivery? And could the Officer Hunt fix the oil leak and subsequently fix the hab? Time was crucial, because we only had limited reserve water bottles and the toilet was also relying on the tank. From personal experience I could tell, holding back on business at the space-loo makes it impossible to stay operational.

The night was awfully cold leading to Sol 4 and so we came up to a new possible reason of the failing water system. Maybe the pipes leading into the hab were frozen? We checked them and although we couldn’t peek inside we could feel how cold they were. The sun still hadn’t turned around far enough to warm up the external water tank. We never had been so anxious for the hot giant star to move faster to hit Mars with a wave of heat.

After two hours of trouble shooting and learning more about our life-sustaining infrastructure we contacted mission control and requested assistance on our problems. We were assured, that we were on the right track and that we will life another day. After we closed the oil leak, which was caused by cap, which wasn’t screwed tight enough, the battery gained quickly power. And once the sun turned the water flooded into our hab.

Because of the crisis situation we had to push our second EVA’s two to three hours back. Me and a crew of three, led by Science Officer Akash Trivedi, were already in the pre-breathing chamber, when we heard the good news of the resolved problems. We could start with a light heart to our mission to collect soil and rock samples.

Akash Trivedi is one of to European members of our crew. The Briton is well connected to the university of Oxford, which asked him to do a so called Matryoshka project. He received satellite data for interesting surfaces on Mars and now wants to collect samples from exactly these spots. Like a Russian doll both elements will complete one another.

For the first time we took the rover out and as we reached our destination climbed on hills to collect the sources. It was fun. The heavy helmet and backpack didn’t really hold me back. But I must say carrying the camera equipment and the necessity to be faster at certain spots and staying longer to have enough flesh for the footage is demanding. It sounds contradicting, but creating great movie footage makes often only sense for the filmmaker.

Coming back was great, because there was a special treat waiting for us. With restored energy and water, we will have our first shower on Mars. We were holding back on it since the beginning, looking on the water crisis, quite a good exercise to deal with such a shortage. So if you excuse me, I have a date with our shower…

Personal Logbook             Today was a fruitful day, but I will be happy when I lay down to sleep.  Partly this is due to our busy schedule, to the high demand of the marsonauts life and the challenges of the work.

Because my crew has little to no experience with documentary work, it is also a workshop for them to align with the demands of filmmaking. In a way, they have to be actors, but actors, who play themselves. Once you understand this and the technical requirements of filmmaking you are half way there.

Some are very talented and pure gold. For others I have to explain certain procedures over and over again. And it’s all good and no ones fault. It just makes it a bit harder for me. Which is tough, when you already do four jobs by yourself.

Today I could also do some more photos on film, which made me very content. During the various problems we had to manage I was doing some photography of our polished helmets and suits. I thought I use the calm to cover theses essentials.

Tomorrow I might stay in the hab for a day. I definitely should take a little rest in one of the next days. I think I will make this decision upon the weather forecast. Still super excited to be here, but taking care of my energy level to continue to deliver good pictures.

Thank you very much for your help and attention.

Willie Schumann, Journalist, Crew 184

Sol Summary – December 4th

MDRS Daily Summary Report for Sol 3

Summary Title: First full up day of activities!

Mission Status: Today really felt like our first day on mars. It was our first full activity day as well as our first day on EVA! Surprisingly it was consensus among our crew that our first full day was actually the easiest day we’ve had so far! I think the excitement of going outside and our team getting into the swing of things has really made us more efficient.

The primary activity for the day was an EVA on all crew to familiarize ourselves with EVA procedures, suit functionality, and operation of the rovers. Many lessons were learned, and it was good that we got the kinks out on a EVA solely suited to it. This will allow our scientific EVAs to be successful later on, as well as ensure that we can safely perform EVAs at a farther distance from the habitat.

Our team is following a specific protocol for depress / repress operations modeled after ISS EVAs. Our team is doing a 25 minute prebreathe in the EVA Prep room to prevent decompression sickness, once complete we are doing a 5 minute depressurization in the actual airlock. During repress we are allocating 5 minutes for repressurization options, though we have the capability to do an ’emergency repress’ if needed at the risk of rupturing crew eardrums.

Below are the several lessons learned we had during EVA.

  1. Peripheral vision is nearly non-existent when wearing the EVA helmets. This makes it difficult to travel together in an ATV convoy as it is impossible to turn around and check on the status of those behind you. We attempted to mitigate this by placing the radios in VOX and keeping in constant voice contact to maintain proper spacing, we also played ‘leapfrog’ with the ATV’s where the leader would stop after several minutes of driving to allow the two behind him to catch up and pass, ensuring no-one was left behind.
  2. We had several communication failures during EVA due to earpieces falling out of ears. We addressed this via using medical tape to secure the earpieces in place.
  3. We had to abort a prebreathe operation due to forgetting to bring supplies into the EVA Prep room. We instituted a supply check prior to closing the Prep Room door. If necessary to break EVA prebreathe protocol we do a 3 to 1 buy back, i.e. if you break prebreathe for 5 minutes, you add 15 minutes to prebreathe to compensate.
  4. Several of our crew members are trained scuba divers. We found that it was difficult to hear other EVA crew while driving the ATV due to engine and wind noise, we instituted a series of hand signals to allow non-verbal communication which quickly and concisely allowed EVA crew status to be conveyed.
  5. Do not wear high pony tails – recommend low braid. A high braid interferes with the way that your head can move within the EVA helmet and vision is confined to roughly a 45 degree angle down and lower.

During the high winds today we heard a bang as the outer airlock door was blown open. The inner airlock door held closed and thus we didn’t all die! We immediately closed the door to the EVA Prep room to form a second seal to the martian atmosphere. During tomorrows familiarization EVA we will enter the EVA Prep room, close the EVA Prep room door, and use that as a depressurization chamber. Immediately upon departure on tomorrows EVA we will repair the outer airlock door.

We also completed our first scientific objectives today, covered below.

Reduced Gravity CPR Experiment: When on Mars there is a possible complication when performing CPR. With your body weight reduced the typical CPR techniques are non functional due to reduced force on the recipient. Our team is practicing techniques under the supervision of our medical officer to allow force application in excess of body weight and measuring output using a springed scaled.

Communication Delay ‘Bricks’ Experiment: Our first two experiment runs of the brick experiment were performed. We started with a 0 time delay baseline, and increased to a 9 minute time delay with verbal information transfer (martian time delay). Our runs will continue to collect more data and determine the most efficient methods of information transfer.

We also completed some major maintenance tasks today

The Musk Observatory has been set up and put through it’s paces and is ready for operation.

Trailer Tank to Static Water Tank transfer operation – Crew completed the transfer this morning completely filling our static water tank. Heartfelt thanks to the Team supporting us for dropping the trailer tank and instructions on how to transfer the water. The MDRS Hab Manual was a roaring success in allowing us to complete this task!

Sol Activity Summary:

  1. Tasks mentioned above
  2. Exercise
  3. Dirt dehydration
  4. Science Dome organization and clean up
  5. Repair of Roof Hatch hole
  6. Our first shower! Heartily enjoying the new heater
  7. Completion of reports (soon)
  1. EVA Request
  2. Sol Summary
  3. Greenhab
  4. Journalist
  5. Operations
  6. Astronomer
  7. Inventory

WiFi range survey Radio check with all radios (all functional except for 1) Look Ahead Plan:

1. Matryoshka EVA #1

2. Communication ‘Lego Bricks’ experiment

3. Sleep Study / crew medical questionnaires

4. Musk Observatory Observation

5. Journalist interviews and filming

6. Medical status surveys with Crew Doctor throughout the day.

7. CPR Experiment

8. HAB Cleaning

9. Radio Range EVA – This activity is cancelled due to 1 EVA per day limit. It will be rescheduled for a future day. We will work with schedulers to ensure they understand this limit for the future.

Anomalies in work:

  1. ATV wobbly wheel
  2. Non Functional Power Strip (removed and replaced)
  3. 3 Non Functional Radios (does not impact operations)
  4. HAB paneling peeling off due to wind damag

Weather: Very mild weather. This was welcome after yesterday.

Crew Physical Status: All crew in good health.

EVA: 2 Familarization EVAs today

Support Requested:

  1. ATV Fuel
  2. Food Drop (not urgent)

.

Astronomy Report – December 4th

Facility Utilized: Musk Observatory

Crewmembers Utilized: Thomas Horn

Supplies Utilized: Solar Telescope / Eyepiece / Camera / Dome

Duration (minutes): 180 minutes

Description of Activity: Radiation is a deadly threat to astronauts both in transit to and on the surface of mars, particularly when the astronauts are exposed to the sun. While astronauts can rely on ground support to identify any solar hazards that develop, communication with the ground cannot be guaranteed. If a Mars Habitat had an independent method of solar observation it could provide critical advance notice that solar flares or other hazards were developing and give direction to crew to shelter in place in radiation storm shelters.

Today’s activity was primarily focused on setting up the solar observatory for operation throughout the mission. A quick survey of the sun was done and no unusual solar activity was spotted. Images below were taken during today’s operation.

Operations Report – December 4th

SOL: SOL 3

Name of person filing report: HuntJ

Non-nominal systems:

  1. Wi-Fi Modem (lack of range).  The wifi in a roughly 5 feet circle from the router, any farther and signal drops off line.  We have tested this and created a signal service map by carrying laptops around the habitat to determine when it picks up signal.  No known fix, will operate internet within this range for now.
  2. HAB Eastern Wall has sustained wind damage, panels are beginning to peel off wall. When wind blows the panels swing and hit support beam causing HAB to vibrate.  In order to fix we need HAB material adhesive delivered, after which we will go EVA and patch the damage,
  3. ATV Yamaha 250 #3 Back Right Wheel is loose, observed to be wobbling during driving.  We will assess a fix during tomorrows EVA – it is possible it simply needs to be tightened, if there is actual mechanical damage then we will move the ATV and declare it off limits.
  4. 3 broken old radios.  All newly delivered radios are functional (thank you supply drop).  Assessing broken radios to determine a fix, if none identified will mark as trash and place in a ‘broken equipment bin’.  4 EVA earpieces were declared nonfunctional, though this may have been due to incorrect settings, will continue to investigate during tomorrows EVA

Generator (hours run): 15.4 Hours

Solar— SOC % : @ 7:34 AM: 65%

                            @ 8:00 AM: 100 % (Generator stopped at this time)

@ 5:34 AM: 42% (Generator started at this time)

Diesel –  48% Full.

Propane –  79 Percent Volume .

Ethanol Free Gasoline (5 Gallon containers for ATV) –   gallons. 2.5 Gallons. One 5 gallon tank empty, second tank half full. Request for extra fuel.

Water (trailer) –  600 Gallons

Water (static) –   500 gallons (We have water again!!)

Trailer to Static Pump used –  Yes

Water (loft) – Static to Loft Pump used –  Yes

Water Meter:  35 gallons

Toilet tank emptied: No

ATV’s Used: 3 Yamaha 250’s and 1 Honda ATV used.

Oil Added? No

ATV Fuel Used:  7.5 gallons used to fill all ATVs to full, though the ATVs did not start with full tanks for our mission.

# Hours the ATVs were Used today: 02:30 hours

Notes on ATVs: See above for wobbly ATV wheel issue.

Deimos rover used: Not used

Hours:

Beginning charge:

Ending charge:

Currently charging:

Sojourner rover used:  ASSIGNED TO DIRECTOR

Hours:

Beginning charge:

Ending charge:

Currently charging:

 

Spirit rover used:  Not used

Hours:

Beginning charge:

Ending charge:

Currently charging:

 

Opportunity rover used:  Not Used

Hours:

Beginning charge:

Ending charge:

Currently charging:

Curiosity rover used:  Not Used

Hours:

Beginning charge:

Ending charge:

Currently charging:

HabCar used and why, where?  No

General notes and comments: Trailer to Static water tank was used today. Pump was dirty and needed cleaning and we did so. Actual process was nominal. We are now full on water again. When inspecting trailer tank water the front tank had a layer of FOD and detritus at the water line. We will keep watch of filter usage in case this increases our burn through rate on filters.

Summary of internet:  Wifi was fixed this morning, though range is limited (see above). Please send hardware manual for the router and modem to allow us to diagnose issues and repair anything that comes up in the future (for example – to allow us to diagnose what certain light combinations signify)

Summary of suits and radios: See above

Summary of Hab operations: Brick Communication Experiment / CPR experiments successful, exercise on all crew. EVA Outer Door repaired.

Summary of GreenHab operations: No fire extinguisher observed – please deliver in future supply drop.

Summary of ScienceDome operations: Soil Salinity experiments (previous crew), dehydrating dirt for use as counterbalance and weights in exercise equipment (filling bags and medicine balls)

Summary of RAMM operations: Not Operational

Summary of health and safety issues: All nominal

Questions, concerns and requests to Mission Support:

Need extra fuel for ATVs. Of the two red fuel cans, one is empty, one is half full.

Pre-Mission food inventory is complete, will be sent as it’s own report. Please reference that report for requested supplies to address gaps in current inventory. Please give ETA for supply drop so we can adequately ration food.

Thank you for your assistance!

GreenHab Report – December 4th

GreenHab Report

Trisha Randazzo

December 4, 2017

Environmental control: (Choose which is appropriate and explain further if needed)

Ambient (no heating or/cooling)

Ambient with window/door open

  • No wind, door open starting at 11:00
  • Plants looked a bit warm so will open door a before 10:00 tomorrow

Heating

  • Functioning nominally

Cooling

  • Not available

Both heating and cooling

  • Nominal

Shade cloth on/off

  • On

Average temperature: (N/A until new sensor is delivered)

  • N/A

Changes to crops: Note all emergence of seedlings, death of seedlings, etc

  • Same tomato is still doing poorly, but watering it
  • Some tomato plants were a bit dropping, possibly due to the heat. Will open Hab door earlier than 11:00

Daily water usage for crops:

  • Complete with no issues

Time(s) of watering for crops:

  • 1:00 PM

Morning research observations: (Currently not operational)

  • N/A

Changes to research plants:

  • N/A

Daily watering and amount of water used:

  • 2 gallons for all soils

Aquaponics: (Currently not operational)

  • N/A

Narrative: Any other information you want to share

  • Paper-whites are looking good!!

Support/supplies needed:

  • Awaiting: